OpenStudy (anonymous):

If $8,000 is placed in an account with an annual interest rate of 4%, how long will it take the amount to quadruple if the interest is compounded annually? Round your answer to two decimal places

OpenStudy (mertsj):

\[32000=8000(1+\frac{.04}{1})^{1(t)}\]

OpenStudy (mertsj):

\[4=1.04^t\] \[\ln 4=t \ln 1.04\] \[t=\frac{\ln 4}{\ln 1.04}=35.3 years\]